OpenStudy (anonymous):

solve the systems of equations by subsitition 4x -3y =3 6x-2y =2

OpenStudy (anonymous):

Y=-1 X= 7/-3

OpenStudy (anonymous):

how did you get that can you show me

OpenStudy (anonymous):

oh wait..

OpenStudy (anonymous):

apologies, x is 0 and y is -1

OpenStudy (anonymous):

can you show me how you got it

OpenStudy (anonymous):

basically multiply the first equation by 6 to get 24X-18Y=18 and then similarly the second equation by 4 to get 24X-8Y=8 then subract the seconds equation from the first; 24X-18Y=18 -(24X-8Y=8) To get; -10Y = 10 Rearannge to get Y=10/-10 hence Y=-1 Subsitute into first equation to get 4X+3=3 hence 4X=0 so x is 0